### v1

# From Tower to Spire: Adding the Speech Modality to a Text-Only LLM

Abstract
Large language models (LLMs) have shown remarkable performance and generalization capabilities across multiple languages and tasks, making them very attractive targets for multimodality integration (e.g., images or speech). In this work, we extend an existing LLM to the speech modality via speech discretization and continued pre-training. In particular, we are interested in multilingual LLMs, such as Tower, as their pre-training setting allows us to treat discretized speech input as an additional translation language. The resulting open-source model, Spire, is able to transcribe and translate English speech input while maintaining Tower’s original performance on translation-related tasks, showcasing that discretized speech input integration as an additional language is feasible during LLM adaptation. We make our code and models available to the community.
From Tower to Spire: Adding the Speech Modality to a Text-Only LLM
Kshitij Ambilduke††thanks: Equal contribution.††thanks: Work begun during an internship at Instituto de Telecomunicações.1, Ben Peters∗2, Sonal Sannigrahi∗2,3, Anil Keshwani†4, Tsz Kin Lam5, Bruno Martins3,6, Marcely Zanon Boito††thanks: Equal contribution.7, André F.T. Martins‡2,3,8,9 1Paris-Saclay University 2Instituto de Telecomunicações 3Instituto Superior Técnico, Universidade de Lisboa 4Sapienza University of Rome 5University of Edinburgh 6INESC-ID 7NAVER LABS Europe 8ELLIS Unit Lisbon 9Unbabel Correspondence: benzurdopeters@gmail.com
1 Introduction
Large language models (LLMs) have demonstrated remarkable success across various text-based natural language processing tasks Achiam et al. (2023); Touvron et al. (2023); Jiang et al. (2024); Yang et al. (2024); Alves et al. (2024); Martins et al. (2024), motivating research into extending them to other modalities. This has led to the development of multimodal LLMs (MLLMs) capable of processing speech, audio, images and video Team et al. (2023); Driess et al. (2023); Rubenstein et al. (2023); Liu et al. (2023); Tang et al. (2023); Défossez et al. (2024); Hu et al. (2024); Laurençon et al. (2024); Huang et al. (2024); Nguyen et al. (2025). For speech-LLM integration, an effortless approach is to link the output of an automatic speech recognition (ASR) system to a text-only LLM Huang et al. (2024). However, this solution fails to fully leverage the LLM’s superior language modeling capabilities for disambiguating transcripts.
More popular are solutions that investigate equipping LLMs with speech processing capabilities through modality projection Shu et al. (2023); Radhakrishnan et al. (2023); Wu et al. (2023a); Tang et al. (2023); Xue et al. (2024); Hu et al. (2024). Typically, a speech foundation model generates speech representations that are mapped to the embedding space of the LLM. The speech model is then fine-tuned along with a projector on speech-to-text tasks to equip the LLM with speech processing capabilities. In this setting, key challenges include prompt overfitting and high training costs, as tuning these MLLMs requires the adaptation of the speech projector module on vast amounts of raw speech data Tang et al. (2023); Hu et al. (2024).
An alternative approach for MLLMs is the use of speech discretization, where continuous speech features are transformed prior to training into sequences of “discrete speech units” (DSUs), which can be processed similarly to text Chou et al. (2023a); Zhang et al. (2023); Rubenstein et al. (2023); Chang et al. (2024); Défossez et al. (2024); Trinh et al. (2024); Maiti et al. (2024); Nguyen et al. (2025). This approach simplifies training by eliminating the need for additional parameters beyond extended embedding matrices. Finally, while both projector-based and discretization-based MLLMs have shown promising results on text-to-speech and/or speech-to-text tasks, their development has prioritized speech-centric tasks at the expense of textual performance. Currently, limited research has focused on integrating speech while preserving the LLM’s original capabilities in textual tasks Chou et al. (2023b); Huang et al. (2024).
In this work we present Spire, a speech-augmented LLM built on top of the open multilingual model Tower Alves et al. (2024). Spire can process English speech and perform ASR and speech translation (ST) while maintaining Tower’s strong performance on machine translation (MT). Spire encodes speech via HuBERT-based Hsu et al. (2021) k-means clusterization, as in previous work Zhang et al. (2023); Rubenstein et al. (2023); Chang et al. (2024). We perform training in two stages: continued pre-training (CPT) and instruction tuning (IT). For the CPT stage, we use a mixture of ASR data and a small fraction of Tower’s text CPT data. For IT, we again leverage Tower’s task-specific MT data, as well as additional English ASR and ST data. Spire is trained using approximately 42.5K hours of speech. Figure 1 illustrates our training process.
We make the following contributions:
We present a pipeline for integrating speech as an additional modality into an existing LLM, enabling it to transcribe and translate English speech while preserving its original MT capabilities;
We compare speech integration at two stages, namely CPT and IT, demonstrating that both stages are essential for achieving optimal performance on speech tasks;
We provide a reproducible pipeline to the community: all our models, datasets and scripts are made available.111https://github.com/utter-project/SpireLM
2 Related Work
Speech-to-Text Models
An increasing number of studies have explored integrating speech into LLMs Zhang et al. (2023); Rubenstein et al. (2023); Hassid et al. (2024). For discrete speech input, Hassid et al. (2024) demonstrate the benefits of initializing a speech LLM from a text-based LLM. SpeechGPT Zhang et al. (2023) applies IT on speech-to-text cross-modal ASR, text-to-speech (TTS), and text-based question answering. AudioPALM Rubenstein et al. (2023) is trained in a multi-task fashion, similarly to SpeechGPT, but on multilingual input. Recently, VoxtLM Maiti et al. (2024) was trained jointly on DSUs and text data for ASR, TTS, and open-ended speech/text generation. Our work is most similar to that of Spirit-LM (Nguyen et al., 2025), which adapts an LLM with an interleaved mixture of DSU and text data which requires an expensive DSU-to-transcript step to create. In contrast, we adopt a more cost-effective input representation that can be extended to any language, regardless of the availability of a speech aligner. Our focus is on successfully incorporating speech input while preserving the original competence of the model, so that the resulting model can successfully perform both speech-to-text and text-only tasks. None of the aforementioned models are trained to preserve the original model’s performance in text tasks.
Adapting LLMs
Previous approaches involve training from scratch with task- and domain-specific data Singhal et al. (2023); Lewkowycz et al. (2022), performing CPT with a diverse training data mix designed to broadly extend the model’s knowledge Wu et al. (2023b), or doing IT use-case-specific data Chen et al. (2023). Recent work has explored combining the latter two approaches Xu et al. (2024a); Alves et al. (2024); Wei et al. (2021); Roziere et al. (2023). In our approach to integrating DSUs into Tower, we take inspiration from Alves et al. (2024) in adopting a two-step CPT+IT process. Our work differs in that we focus on adding the speech modality, whereas Alves et al. (2024) focused on increasing the multilingual capabilities of an LLM.
Continuous and Discrete Speech Representations
Self-supervised speech representation models produce contextualized high-dimensional speech vectors directly from raw audio Hsu et al. (2021); Baevski et al. (2020); Chen et al. (2022), largely outperforming statistical speech features on downstream tasks Yang et al. (2021). These continuous representations can be used to derive DSUs that capture both linguistic content and prosody through clustering Borsos et al. (2023); Kharitonov et al. (2022). DSUs provide better alignment with textual data, facilitating the transfer of successful training settings from the text domain Cui et al. (2024). Building on Lakhotia et al. (2021), which demonstrated that HuBERT Hsu et al. (2021) is a powerful feature extractor, several studies have adopted this approach, incorporating a k-means clustering step for discretization Zhang et al. (2023); Rubenstein et al. (2023); Lam et al. (2024); Chang et al. (2024); Nguyen et al. (2025). Xu et al. (2024b) study the optimal settings to obtain DSUs in terms of cluster size and feature extraction layer. We use their findings to inform our initial choices.
3 Spire: A Speech-to-Text LLM
Our goal is to equip an LLM with speech capabilities while preserving its preexisting text capabilities. As our starting point, we select Tower Alves et al. (2024) as our LLM, which was developed from Llama-2 Touvron et al. (2023) with a two-step approach: CPT on a mixture of monolingual and parallel data (TowerBase), followed by IT on translation-related tasks (TowerInstruct). We use a similar approach to extend Tower to speech. First, we perform CPT with a combination of text-only and aligned speech-to-text datasets, followed by IT using both text-only general-purpose and task-specific data curated in TowerBlocks,222https://huggingface.co/datasets/Unbabel/TowerBlocks-v0.2 alongside task-specific speech-to-text datasets. We name our model Spire.
3.1 Speech Discretization
To more easily transfer the training set-up of Tower, we use DSUs as opposed to an auxiliary speech encoder. For all speech datasets that were used, we follow recent discretization methodology Zhang et al. (2023); Rubenstein et al. (2023); Chang et al. (2024) to produce DSUs by first extracting continuous speech representations for our speech data from the 22nd layer of an English HuBERT-large model, and then using k-means clustering () to produce centroids that are used to convert our continuous speech representation into a discrete sequence of cluster IDs.333Optimizing the layer selection for feature extraction is a complex research problem Pasad et al. (2023); Mousavi et al. (2024). In this work we follow the insights from Gow-Smith et al. (2023) and Xu et al. (2024b). We train our k-means model on a collection of 235K audio files (approximately 720 hours), drawn from three speech corpora: CoVoST-2 (Wang et al., 2021b), VoxPopuli Wang et al. (2021a), and Multilingual Librispeech (MLS; Pratap et al., 2020). The CoVoST subset consists of 62K audio files from 10,049 speakers, with a maximum of 8 audio files per speaker. The VoxPopuli subset includes 65K audio files from 639 speakers, capped at 250 audio files per speaker. Finally, the MLS subset contains 107K audio files from 5,490 speakers.
3.2 SpireBase
SpireBase is trained from TowerBase-7B using both text-only and aligned speech-to-text datasets. Following previous work, we incorporate a fraction of Tower’s original training data to preserve its original performance Scialom et al. (2022); de Masson D’Autume et al. (2019).
3.2.1 Data
We use a mixture of monolingual and parallel text in Chinese (zh), Dutch (nl), English (en), French (fr), German (de), Italian (it), Korean (ko), Portuguese (pt), Russian (ru), and Spanish (es), that was sourced from the Tower training data, as well as English ASR data sourced from popular open-source ASR datasets, as reported in Table 2. Both speech and text data are downsampled to create a 6B token data mixture (5B speech; 1B text), measured by the model tokenizer.444Preliminary experiments on the data mixture led to this particular choice. Note that the 5B speech tokens include both DSUs (4.4B tokens) and their text transcriptions (0.6B tokens).
Text Data
The monolingual text data split corresponds to data from mC4 Raffel et al. (2019), a multilingual web-crawled corpus which we uniformly sample from across all languages. The parallel data split includes uniformly sampled instances to and from English (enxx) for the 10 languages, sourced from various public sources. Further details can be found in Alves et al. (2024).
Speech Data
3.2.2 CPT Setup
We train SpireBase using the MegatronLLM codebase Cano et al. (2023) on 8 A100-80GB GPUs for 6 days. We use the same hyperparameters as Tower, except for the effective batch size, which in our case is 2,304. To incorporate the DSUs in the CPT stage, we extend the model’s original vocabulary by 5000 types, e.g., <extra_id_x>. This allows us to have a vocabulary that can encode both text in subword units and speech in DSUs. For the extended vocabulary, we initialize new embeddings from a multivariate Gaussian distribution. The mean of this distribution is set to the average of the original embeddings, while the covariance is derived from the empirical covariance of the original embeddings, scaled by a factor of Hewitt (2021).
3.3 SpireFull
The SpireFull model is obtained by performing instruction tuning on SpireBase using task-specific text-only and aligned speech-to-text data.
3.3.1 Data
We use a mixture of text and speech instructions for ASR, MT, and ST. The prompt formats used during training are shown in Table 1.
Text Data
We use TowerBlocks Alves et al. (2024), which includes high quality translation bi-texts between English and the other languages supported by Tower. It also includes instructions for the translation-related tasks of named entity recognition and automatic post-editing.
ASR Data
ST Data
In our IT set, we use 842 hours of speech across three ST training sets: FLEURS (all nine language pairs; we filter out examples whose transcriptions overlap with the FLORES devtest set), Europarl-ST Iranzo-Sánchez et al. (2020) (en {de, es, fr, it, nl, pt}), and CoVoST-2 (enzh). Since this amounts to far less data for ST than what is available for ASR, and since en{ko, ru} have only examples from the tiny FLEURS set, we augment our speech collection with pseudo-labeled data. This has been shown to be effective for other ST systems (Barrault et al., 2023). For the CV, SPGI, and GigaSpeech datasets, we select 300K ASR samples to be pseudo-labeled each. These examples are translated to all nine target languages using TowerInstruct-13B.555https://huggingface.co/Unbabel/TowerInstruct-13B-v0.1 Although this produces a very large ST corpus, not all predictions are of high quality, so we filter out examples for which the transcript-translation combination has a COMET-QE666https://huggingface.co/Unbabel/wmt22-cometkiwi-da (Rei et al., 2022b) score under 85. Finally, for each language pair, we sample 60K examples to be used in direct ST prompts and another 60K samples to be used in multi-turn prompts. This process results in 180K direct ST prompts and 180K multi-turn prompts for each language pair.777Due to our aggressive filtering, we were left with slightly fewer examples for en zh.
| ASR (CPT) | |||
| MT (CPT) | |||
| ASR (IT) | |||
| Direct ST (IT) | |||
| Multi-turn ST (IT) | |||
| Dataset | Task | Phase | # DSUs | # Hours |
| SPGI Speech | ASR | CPT | 645M | 5.1K |
| Gigaspeech | ASR | CPT | 1.2B | 9.9K |
| MLS | ASR | CPT | 2.4B | 19.2K |
| VoxPopuli | ASR | CPT | 69M | 0.5K |
| CV | ASR | IT | 105M | 0.8K |
| Europarl-ST | ST | IT | 122M | 1.0K |
| FLEURS | ST | IT | 11M | 0.09K |
| CoVoST-2 | ST | IT | 12M | 0.09K |
| SPGI Speech | Pseudo-ST | IT | 350M | 2.8K |
| GigaSpeech | Pseudo-ST | IT | 161M | 1.3K |
| CV | Pseudo-ST | IT | 212M | 1.7K |
3.3.2 IT Training Setup
Similar to Tower, we use the chatml template OpenAI (2023) to format our instructions in dialogue form. We train models using Axolotl888https://github.com/axolotl-ai-cloud/axolotl on 4 H100-80GB GPUs for 2.7 days. We use a learning rate of and a cosine scheduler with 100 warm-up steps. We train for 4 epochs with an effective batch size of 576 and a weight decay of 0.01. We impose a maximum sequence length of 4096 and use the AdamW optimizer Loshchilov and Hutter (2019). Other hyperparameters are derived from TowerInstruct (Alves et al., 2024).
4 Experiments
We evaluate our models across three tasks: ASR, MT, and ST. First, we present our results for ASR (§4.1), confirming the competitive performance of Spire in the speech domain. We then present MT results (§4.2), demonstrating that the speech performance does not come at the expense of the original model’s MT performance. Finally, we present results for ST (§4.3) to investigate model performance on a task that requires both ASR and MT capabilities.
Our Models
Across experiments, we compare all models shown in Table 4. These models vary in which base model they were trained on, whether they underwent speech-centric CPT, and on what set of instructions IT was performed. Apart from Spire models, we report a few ablation studies in the IT stage where:
i) no CPT was performed (TowerFull);
ii) no data from TowerBlocks was seen during IT (SpireNoBlocks), and
iii) pseudo-labeled ST data and FLEURS were omitted (SpireNoPseudo).
When reporting results, we highlight toplines when possible.
| Model | Base Model | CPT | IT | |||
| Speech | Text | Speech | Pseudo | Text | ||
| TowerFull | TowerBase-7B | ✗ | ✗ | ✓ | ✓ | ✓ |
| SpireBase | SpireBase | ✓ | ✓ | ✗ | ✗ | ✗ |
| SpireFull | SpireBase | ✓ | ✓ | ✓ | ✓ | ✓ |
[…]
| SpireNoPseudo | SpireBase | ✓ | ✓ | ✓ | ✗ | ✓ |
Evaluation Setup
Across models and tasks, we perform inference with greedy decoding with a maximum of 256 generated tokens. For the Tower and Spire models, we decode with vllm. However, since vllm does not support all of our baselines, we use alternative libraries (namely, transformers) where necessary. Unless specified otherwise, we use zero-shot prompts for all models and tasks.
4.1 ASR Experiments
Datasets and Metrics
We evaluate ASR performance across multiple test sets, in order to cover a variety of recording styles: Librispeech (LS) test-clean and test-other Panayotov et al. (2015), FLEURS Conneau et al. (2023), and VoxPopuli.999For CPT models, LS is an in-domain evaluation because its training set is part of MLS. We report the Word Error Rate (WER) between the hypotheses and gold transcripts, after Whisper normalization (Radford et al., 2023).
Baselines
We report results for the following models:
Whisper (Radford et al., 2023) is an encoder-decoder transformer model trained on over 5 million hours of labeled data, it performs multilingual ASR and to-English ST. We report results for the 74M parameter Whisper-base and the 1.5B parameter Whisper-large-v3 versions.
SeamlessM4T (Barrault et al., 2023) is an encoder-decoder transformer trained on 406K hours of speech that performs ASR, ST and MT across 100 languages. We report results for the 2.3B parameter SeamlessM4T-large-v2 version of the model.
Spirit-LM Nguyen et al. (2025) is the most similar work to ours. It is a decoder-only model, trained from Llama-2 on 307B tokens of text, 458K hours of unlabeled speech, and 111K hours of labeled speech. Unfortunately, despite the availability of their inference code, we were unable to reproduce its reported performance on speech tasks.
HuBERT-large+CTC is a CTC-based ASR model trained using the same speech representation model we use for DSU generation, and using the same ASR data from the IT stage (Section 3.3.1). This model allows us to compare our IT-only TowerFull model against a model which has access to continuous speech representations.101010The hyperparameters for this ASR model are described in Appendix B.
| LibriSpeech | FLEURS | VoxPopuli | ||
| Clean | Other | |||
| Baselines | ||||
[…]
| HuBERT-large+CTC | 4.3 | 7.6 | 11.4 | 14.7 |
| Spirit-LM | 6.0* | 11.0* | - | - |
| SeamlessM4T | 2.6 | 4.9 | 8.1 | 7.5 |
| Whisper-base | 5.0 | 11.9 | 12.1 | 9.8 |
| Whisper-large-v3 | 1.8 | 3.7 | 5.8 | 9.2 |
| Our models | ||||
[…]
| TowerFull | 9.5 | 13.8 | 14.3 | 40.7 |
| SpireNoBlocks | 4.1 | 7.4 | 10.4 | 15.8 |
| SpireNoPseudo | 3.9 | 7.3 | 11.1 | 14.3 |
| SpireBase | 28.9 | 56.3 | 11.0 | 13.7 |
| SpireFull | 4.2 | 7.1 | 10.7 | 15.8 |
*We were unable to reproduce Spirit-LM’s ASR performance; therefore, we report their self-reported LS results using ten-shot prompts
Results
Our results are presented in Table 4. SpireFull’s performance demonstrates that performing both the CPT and IT stages is an effective strategy to give speech capabilities to a text LLM. Notably, SpireFull outperforms the HuBERT-large+CTC baseline on three out of four datasets—an impressive result given that the CTC model has a helpful inductive bias and access to continuous features, both of which SpireFull lacks.
The performance gap between SpireFull and TowerFull (5.3 points in LS test-clean) demonstrates that combining CPT and IT is more effective than using IT alone. We further observe that while TowerFull obtains better results than SpireBase, it performs worse than HuBERT-large+CTC, showing that the CPT stage is crucial in outperforming a model that has access to continuous features. Additionally, the minimal difference between SpireNoBlocks and SpireFull in the IT stage suggests that incorporating textual tasks does not negatively impact ASR performance. For SpireBase, it is surprising that FLEURS and VoxPopuli results were somewhat strong zero-shot settings, given that non-instruction-tuned models often struggle to work in out-of-domain without in-context learning examples.111111We also tried prompting SpireBase with few-shot examples, but the results were significantly worse. This may be due to the length of the DSU sequences, which likely led to in-context examples that were too long for the model to handle effectively.
Finally, although SpireFull cannot match the performance of SeamlessM4T or Whisper-large-v3, both of which were trained on far more speech data, it does exceed the performance of Whisper-base on both sections of LS and FLEURS. It also outperforms Spirit-LM on LS, which is notable because both models are derived from Llama-2 and make use of HuBERT DSU tokens.
4.2 MT Experiments
| enxx | xxen | |||
| C22 | spB | C22 | spB | |
| Baselines | ||||
[…]
| SeamlessM4T | 87.22 | 39.0 | 87.42 | 39.9 |
| Towerbase-7b | 87.38 | 37.8 | 88.02 | 41.7 |
| Towerinstruct-7b | 88.45 | 38.8 | 88.27 | 42.0 |
| Our models | ||||
[…]
| TowerFull | 88.57 | 39.4 | 88.17 | 41.7 |
| SpireNoBlocks | 82.98 | 34.2 | 85.93 | 36.1 |
| SpireNoPseudo | 88.40 | 38.9 | 88.22 | 42.0 |
| SpireBase | 87.41 | 37.4 | 87.97 | 41.4 |
| SpireFull | 88.54 | 39.3 | 88.21 | 41.8 |
| ende | enru | enzh | deen | ruen | zhen | |||||||
| C22 | spB | C22 | spB | C22 | spB | C22 | spB | C22 | spB | C22 | spB | |
| Baselines | ||||||||||||
[…]
| SeamlessM4T | 77.76 | 27.8 | 83.22 | 34.2 | 80.14 | 29.7 | 78.69 | 26.6 | 80.58 | 32.5 | 76.96 | 23.8 |
| Towerbase-7b | 79.96 | 36.1 | 83.08 | 34.2 | 83.49 | 33.3 | 83.56 | 41.1 | 80.06 | 32.7 | 78.48 | 23.5 |
| Towerinstruct-7b | 82.34 | 38.8 | 84.66 | 34.9 | 85.09 | 35.3 | 84.95 | 45.1 | 82.94 | 36.7 | 80.14 | 26.1 |
| Our models | ||||||||||||
[…]
| TowerFull | 82.63 | 39.2 | 84.55 | 34.5 | 85.39 | 37.2 | 84.65 | 45.2 | 82.41 | 35.6 | 79.68 | 25.9 |
| SpireNoBlocks | 67.97 | 24.4 | 73.86 | 26.6 | 77.80 | 29.6 | 73.24 | 28.7 | 78.09 | 29.1 | 73.01 | 17.6 |
| SpireNoPseudo | 82.18 | 38.5 | 84.31 | 34.7 | 85.31 | 37.6 | 85.04 | 45.5 | 82.56 | 36.2 | 79.91 | 26.2 |
| SpireBase | 79.88 | 34.7 | 83.04 | 33.7 | 83.85 | 32.4 | 83.19 | 40.5 | 80.20 | 32.4 | 78.65 | 23.1 |
| SpireFull | 82.50 | 39.5 | 84.60 | 34.9 | 85.37 | 37.3 | 85.24 | 45.2 | 82.58 | 36.4 | 79.92 | 26.3 |
Having demonstrated that our CPT and IT approaches work well for ASR, we now turn to MT. The key question is whether Spire can maintain Tower’s strong performance on MT, despite its speech-centric CPT. We report performance for translation-related tasks in Appendix C.
Datasets and Metrics
We use two datasets for MT: FLORES-200 Team et al. (2024), which covers all of our models’ languages, and the WMT23 test set (Kocmi et al., 2023), which covers en{de, ru, zh}. We report COMET-22 Rei et al. (2022a) and spBLEU121212nrefs:1|case:mixed|eff:no|tok:flores200| smooth:exp|version:2.5.1 Papineni et al. (2002) scores via the SacreBLEU toolkit Post (2018).
Baselines
We compare the Spire models against the text-to-text translation performance of SeamlessM4T. Additionally, we report the performance of Towerbase-7b and Towerinstruct-7b as toplines.
Results
Our results showcase that even after the speech-centric CPT and mixed speech and text IT stage, the Spire models retain Tower’s performance on both FLORES (Table 5) and WMT23 (Table 6). This indicates that neither CPT nor IT on speech data negatively impact the model’s ability to perform MT. This is true for both CPT-only models, where SpireBase achieves performance comparable to Towerbase on both datasets; and for IT models, where SpireFull and TowerFull both perform slightly better than Towerinstruct on enxx, which is possibly an artifact of the large number of multi-turn enxx ST instructions in their IT set. Notably, our strongest Spire model also outperforms SeamlessM4T by both metrics on all WMT23 language pairs, and for both enxx and xxen on FLORES.
4.3 ST Experiments
| FLEURS | CoVoST-2 | |||||||||||
| Direct | Self-Cascade | Gold | Direct | Self-Cascade | Gold | |||||||
| C22 | spB | C22 | spB | C22 | spB | C22 | spB | C22 | spB | C22 | spB | |
| Baselines | ||||||||||||
[…]
| SeamlessM4T | 84.63 | 33.7 | 75.45 | 22.4 | 86.79 | 38.7 | 84.79 | 36.8 | 72.36 | 19.4 | 86.55 | 39.0 |
| Our Models | ||||||||||||
[…]
| TowerFull | 79.10 | 26.1 | 83.42 | 31.9 | 88.27 | 39.2 | 71.52 | 20.1 | 74.17 | 25.8 | 87.14 | 38.5 |
| SpireNoBlocks | 81.11 | 27.1 | 79.46 | 28.9 | 82.44 | 34.0 | 74.02 | 23.2 | 68.09 | 22.8 | 69.31 | 26.8 |
| SpireNoPseudo | 62.80 | 11.7 | 83.79 | 32.2 | 88.10 | 38.7 | 59.88 | 6.8 | 78.15 | 29.7 | 87.10 | 39.0 |
| SpireFull | 81.33 | 27.1 | 85.21 | 33.7 | 88.36 | 39.2 | 74.25 | 23.2 | 78.78 | 30.0 | 87.11 | 38.5 |
As Spire has shown success at both ASR and MT, we now investigate its performance on ST.
Datasets
ST approaches
We evaluate ST performance on both direct ST and self-cascaded ST,131313We also tried inference with the multi-turn prompt format shown in Table 1, but results were similar to a self-cascade. Reporting the self-cascade enables comparison to models that do not support the multi-turn format, i.e. SeamlessM4T. in which each model transcribes the audio before translating its own output to the target language (i.e., ASR followed by MT). To assess the impact of ASR error propagation, we also report MT results given gold transcriptions.
Results
Our ST results on FLEURS and CoVoST-2 are presented in Table 7. Among our models, the ones that were trained on large quantities of pseudo-labeled ST data (TowerFull, SpireNoBlocks, and SpireFull) achieve far higher scores on direct ST than the one that did not (SpireNoPseudo). This indicates that merely performing CPT with ASR and MT data is not enough to achieve generalization to the task of direct ST, even if the model excels at both ASR and MT. Indeed, we also attempted direct ST with SpireBase and it failed to produce output in the target language, even when given few-shot prompts.
We also observe that SpireNoBlocks performs nearly as well at direct ST as SpireFull, even though its MT performance is much poorer (see Table 5 and 6), showing that, surprisingly, competence at MT is not very helpful for direct ST. SpireFull achieves the best self-cascaded performance by a significant margin for both datasets, outperforming both models that were trained with fewer speech samples (TowerFull and SpireNoPseudo), and the model that was not tuned on MT (SpireNoBlocks). This suggests that, unlike direct ST, both ASR and MT competencies are necessary for a strong self-cascade performance. Although SpireFull does not reach the direct ST performance of SeamlessM4T, it manages to achieve competitive performance despite using far less ST data Barrault et al. (2023). Our self-cascading experiments additionally demonstrate that SpireFull maintains greater robustness to its own outputs than SeamlessM4T.
5 Conclusion
In this work we presented Spire, a simple and effective recipe for adapting a text-based, translation-specialist LLM to the speech modality while preserving the original performance on text-based tasks. We investigated the impact of speech integration on two stages of LLM adaptation, CPT and IT, finding that both contribute to the final model’s performance on speech tasks. Our results demonstrate that we are able to successfully integrate a new modality without compromising the original model’s capabilities. Spire achieves competitive performance on ASR, while its MT abilities remain on par with the original Tower model. Finally, for the ST task, we find that the leveraging ASR and MT data does not directly transfer to ST performance. Nonetheless, the model achieves promising performance with both direct and self-cascaded ST.
As future work, we intend to extend this recipe to multilingual settings by replacing our English HuBERT speech component by the multilingual mHuBERT-147 Boito et al. (2024). We also plan to leverage the flexibility of DSU modeling to investigate the integration of speech generation tasks. To benefit the research community, we only use publicly available and licensed data to train our models, making our results reproducible.
Limitations
The downstream tasks we evaluate on are restricted to MT and ASR/ST, which provide an idea of the model performance but do not give us the full picture. We plan to address this by utilizing the LM-harness evaluation Gao et al. (2024) to evaluate on a suite of text-based benchmarks such as MMLU (Multitask Language Understanding) Hendrycks et al. (2021b, a), Arc (Commonsense Reasoning) Clark et al. (2018), Belebele (Reading Comprehension) Bandarkar et al. (2024), and HellaSwag (Sentence Completion) Zellers et al. (2019). Lastly, our model handles speech and text on the input side but is currently limited to generating only text.
Acknowledgments
This work was supported by EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), by UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee (grant number 10039436: UTTER), by the project DECOLLAGE (ERC-2022-CoG 101088763), by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Center for Responsible AI), by Fundação para a Ciência e Tecnologia (FCT) through the project with reference UIDB/50021/2020 (DOI:10.54499/UIDB/50021/2020), and by FCT/MECI through national funds and when applicable co-funded EU funds under UID/50008: Instituto de Telecomunicações. This work was performed using HPC resources from GENCI–IDRIS (Grant 2023-AD011014668R1). We thank Duarte Alves and Giuseppe Attanasio for their insightful comments.
Appendix A Data
A.1 Speech Data Preprocessing
Normalization
In order to make transcripts consistent across the different datasets, the following normalization is applied:
GigaSpeech (CPT): we lower-case the text and replace punctuation tags: <COMMA>, <PERIOD>, QUESTIONMARK>, <EXCLAMATIONPOINT> with their appropriate punctuation.
MLS (CPT): we apply a tail-end normalization step here which uniformly samples each speaker to have at maximum 13 transcriptions. This allows us to have a better distribution of speakers.
CV (IT): we subsampled from CommonVoice to ensure a minimum duration of 3 seconds per sample. To enhance transcript diversity, we limit each transcript to 4 unique speakers.
Deduplication
Appendix B CTC-based ASR model
We train a CTC-based ASR model using the HuggingFace Transformers library Wolf (2019), leveraging the ASR data from the IT stage (CommonVoice, Table 2) as training data. Our ASR model is made of the HuBERT-Large141414https://huggingface.co/facebook/hubert-large-ll60k speech representation model, followed by three hidden layers and a vocabulary projection layer. We train for 50 epochs with a dropout of 0.3 and a learning rate of 1e-4 with a warm-up ratio of 0.15. The best checkpoint is selected using CER scores. This was obtained at step 220K (at epoch 12.8).
Appendix C Translation-related Tasks
Tower was evaluated on translation-related tasks in addition to MT. We follow their evaluation setup and use the Tower-eval suite151515https://github.com/deep-spin/tower-eval Alves et al. (2024) to evaluate Spire models on APE and NER. For a detailed description of the tasks, we refer the readers to Alves et al. (2024). Briefly, APE measures final translation quality on WMT23 after post-editing with NLLB-3.3B Team et al. (2022), and NER measures entity recognition on MultiCoNER 2023 Fetahu et al. (2023) test. We report COMET-22 for APE and sequence F1 score for NER. We evaluate APE for en{de, ru, zh} and NER for {de, en, es, fr, it, pt, zh}.
Table 8 reports our results. We achieve comparable scores to Towerinstruct across both tasks and all language directions considered. Notably, we outperform TowerFull in all settings, which may be suggestive of the benefit of including text data in the CPT stage.
| APE | NER | ||
| enxx | xxen | Multilingual | |
| TowerInstruct-7b | 83.08 | 80.29 | 71.56 |
| TowerFull | 82.65 | 79.90 | 65.07 |
| SpireFull | 83.13 | 80.08 | 67.10 |